OpenStudy (anonymous):

If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom worked together to complete it? (A) 0.8 days (B) 1.09 days (C) 1.23 days (D) 1.65 days

hero (hero):

Let s = Number of Days Sam Does A Job l = Number of Days Lisa Does A Job t = Number of Days Tom Does A Job Then: \[\frac{slt}{sl + st + lt}\]

hero (hero):

In other words, Number of days then do it together is: \[\frac{4 \dot\ 6 \dot\ 2}{(4)(6) + (4)(2) + (6)(2)}\]

OpenStudy (raden):

v1+v2+v3 = x/4 + x/6 + x/2 v1+v2+v3 = (3x+2x+6x)/12 v1+v2+v3 = 11x/12 v1+v2+v3 = x/(12/11) so, if they are do it together for 12/11 = 1.09

OpenStudy (raden):

just same like @Hero

hero (hero):

Mines is simpler bro

OpenStudy (raden):

yea, i see...